There are five envelopes in a box. One envelope contains a penny, one a nickel, one a dime, one a quarter, and one a half-dollar. A person selects an envelope. Find the expected value of the draw

Answer by stanbon(75887)

There are five envelopes in a box. One envelope contains a penny, one a nickel, one a dime, one a quarter, and one a half-dollar. A person selects an envelope. Find the expected value of the draw
------
E(x) = (1/5)(1+10+25+50) = (1/5)(86) = 17 1/5
-------------------
Cheers,
Stan H